OpenStudy (anonymous):

Start with 2 rabbits. Each pair has 4 babies per month. What's the equation ? I already figured out how many rabbits are in each month. I just need an equation.

OpenStudy (anonymous):

\[2*4^{n}=rabbits\]

OpenStudy (anonymous):

Month 0= 2 rabbits Month 1= 6 Month 2=18 Month 3= 54

OpenStudy (anonymous):

That doesn't work

OpenStudy (paxpolaris):

divide your numbers by 2 ... that might make the answer more obvious

OpenStudy (anonymous):

It is not a plus its a multiplicacion: \[2*4^0=2\] \[2*4^1=8\] \[2*4^2=32\] And so on...

OpenStudy (anonymous):

Where are you getting the 8 from? It's 6 rabbits for month 2

OpenStudy (anonymous):

I thought it was exponential

OpenStudy (paxpolaris):

Month 0: 2 rabbits = 1 pair Month 1: 6 = 3 pair Month 2:18 = 9 pair Month 3= 54 = 27 pair is the pattern more obvious now ...

OpenStudy (anonymous):

I get the pattern. I need a equation

OpenStudy (paxpolaris):

2*3^x

OpenStudy (anonymous):

THAT WORKS. YOU ARE SO SMART. YOU RE SO SMART OH MY GOD. THANK YOU. IM CRYING. (I get emotional okay)

OpenStudy (anonymous):

that works for almost everyone of them except month 0.

OpenStudy (anonymous):

is it okay if it doesn't work for month 0??

OpenStudy (anonymous):

Month \[2*3^0=2*1=2\]

OpenStudy (anonymous):

Oh yeah. Totally forgot it equals 1. Thank you! You two are my heroes.